When corn sells for $4 a bushel and fuel sells for 42 a gallon: What is the equation for the supply curve?

The supply of wheat is given by the following equation: Qws = -6 + 4Pw - 2Pc - P, where Qws is the quantity o wheat supplied, in millions of bushes; Pw is the price of wheat per bushel; Pc is the price of corn per bushel; and Pf is the price of tractor fuel per gallon
Answer #1

Answer

Supply of wheat is given by the following equation: Qws = -6 + 4Pw - 2Pc - P-----------------------(1)

Supply curve shows the relation between Quantity supplied and Price of a good.

Here Good is wheat and hence supply curve shows the relation between Quantity supplied of wheat(Qws) and Price of wheat(Pw)

It is given that corn sells for $4 a bushel and fuel sells for 42 a gallon => Pc = 4 and P = 42

Putting this in (1) we get :

Qws = -6 + 4Pw - 2Pc - P

=> Qws = -6 + 4Pw - 2*4 - 42

=> Qws = 4Pw - 56 --------------------- Supply equation.

Note :

When you draw Supply curve you put Q on horizontal axis and P on Vertical axis and because of this in order to draw a supply curve we write Supply as P as a function of Q.

Here we write this as Pw = (1/4)(Qws + 56)

=> Pw = 14 + 0.25Qws. --------------------------Inverse supply equation

This is what we called Inverse supply(where we write P as a function of Q).